Tevatron Higgs Searches in Run II

Pekka Sinervo University of Toronto

Overview

• Overview of Higgs signatures

• Higgs Production and Decay

• Tevatron and CDF/D0 / upgrades

• B Tagging Capabilities and Jet Resolution • Sensitivity for a Standard Model Higgs • SUSY Higgs detection

• Summary

The Standard Model

The ingredients of the SM are

• 12 fermions – 6 quarks (u, d, s, c, b, t)

– 6 leptons (e^{−}, ν_{e}, μ, ν_{μ}, τ, ν_{τ})

• 3 electroweak bosons (W^{+} , W^{−}, Z^{∘}, γ)

• 8 gluons to describe the strong force

• ?? to describe gravity

Enormously successful paradigm since ∼ 1975

• Predicted rich phenomenology

• Accurate! (eg. δ_{M_{z}}/M_{z} ∼ 8 × 10^{−5})

• Successfully predicted existence of charm and top quark

Higgs boson the only “essential” missing element

• Represents a significant “vacuum” of knowledge

• Its discovery has the highest scientific urgency

Brief Overview of Higgs Phenomenology

Higgs allows “spontaneous symmetry breaking” in SM

• Sets the EWK mass scale

• Yukawa interactions with fermions generates mass

It couples to quarks, leptons and EWK gauge bosons

• Production rate set by mass & couplings to “beam”

• Decays are mediated by same couplings

• In p p collisions, produced

– via qq̄ annihilations through ∗ virtual W/Z for light quark initial states

∗ directly via bb̄ initial states

– indirectly in gg interactions through gg → t t → H

Current Limits on the Higgs

Data point to a low mass Higgs

• LEP II data excludes M_{H} < 113 GeV/c^{2} – suggests an effect around 115 GeV/c^{2}

• Tevatron & LEP II data imply M_{H} < ∼ 200 GeV/c^{2}

The Higgs sector is already constrained

• LEP II limit of 113 and ∼ 200 GeV/c^{2}

• A Higgs discovery could be just around the “corner”

Features of a Light SM Higgs

Higgs couplings to fermions & EWK bosons specify BR

• In region M_{H} < 130 GeV/c^{2}, H → bb̄ dominates

• For M_{H} > 130 GeV/c^{2}, H → W W^{∗} takes over

• Cross over is ∼ 135 GeV/c^{2}

• Other decay modes important, but are not relevant to a Tevatron search

Dicus, Steltzer, Sullivan & Willenbrock, PRD D59 (1999) Spira, private comm.

Higgs Production Cross Section

Cross section at Tevatron energies miniscule

• Arises from small couplings to light fermions • Makes “associated” production quite important

Some relevant points of reference:

• Total cross section ∼ 7 × 10^{10} pb

• Smallest measured total cross section ∼ 5 pb • Lowest upper limits ∼ 0.1 − 0.2 pb (W′)

Tevatron Upgrades for Run II

Tevatron has been significantly improved for Run II

• Main Injector replaces Tevatron for injection – Better p̄ transfer efficiency

– Emittance better-matched at injection

• Antiproton Recycler – effective doubling in p̄ production rate • Increase in √s from 1.8 to 2.0 TeV

⇒ 30-40% increase in production σ

• Go from 6 × 6 → 36 × 36 operation – Reduces interactions per crossing

Performance goals for collider are

• Increase typical L_{max} from 1.6 × 10^{31} to 2 × 10^{32} cm^{−2}s^{−1} • Deliver ∼ 2 fb^{−1} by 2002 (Run IIa)

• If successful in increasing p̄ availability – Go to 108 × 108 operation

– Increase L_{max} to 5 × 10^{32} cm^{−2}s^{−1}

– Deliver ∼ 15 fb^{−1} by end of 2007 (Run IIb)

CDF Upgrades for Run II

CDF has undergone major upgrades for Run II

• Tile-scintillating fibre calorimeters in 1.0 < |η| < 3.6 • Central outer tracker (COT)

• SVX II and ISL silicon tracking systems

• Fully pipelined electronics and new L1/L2/L3 triggers • Upgraded muon systems in 0.8 ≤ |η| ≤ 1.2

• C++ offline reconstruction and new processing farm

D0 / Upgrades for Run II

D0 / has undergone a similar transformation

• New charged particle tracking system – 2.0 T solenoid

– Inner silicon microstrip tracker (SMT)

– New scintillating fibre tracker (CFT)

• Preshower detectors to improve e/π separation

• Forward muon system chambers & scintillators

• Restructured L1/L2/L3 trigger, electronics and DAQ SUSY Higgs Workshop, Mar 2001

Higgs Detection during Run II

SM Higgs production is in principle detectable

• Search of Run I data is clear demonstration – see G. Watts’ parallel talk

• For case of M_{H} = 120 GeV/c^{2}, cross section – for gg → H_{SM} is ∼ 0.9 pb

– for qq̄′ → W H_{SM} is ∼ 0.2 pb

– for qq̄ → Z H_{SM} is ∼ 0.1 pb

• Issues for detection are:

– Reconstruction technique and signal efficiency – Background rejection

Run II SUSY/Higgs Workshop has focussed activity

• Summarizes multiple studies done on – SM Higgs production

– MSSM Higgs scenarios

• Results provided in terms of ∫ Ldt

• Report from workshop now available (hep-ph/0010338)

Workshop Tools

Participants explored complementary approaches

1. Used existing CDF results and extrapolated • Accounted for SVX improvements

• Took into account new triggers

• Extrapolated jet resolution improvements

2. Developed independent detector simulation (“SHW”) • Parametrized resolutions and acceptances

• Took account of D0 / and CDF upgrades – Provide model of “average” detector

3. Assessed neural-net and “advanced” analysis techniques

Results provided for many scenarios and channels

• Resulted in a fairly coherent set of predictions – Assessed impact of “combined” statistical power

– Also noted some areas where uncertainties remain large

• Signal-to-noise issues are key to success

B Tagging Issues

“B Tagging” an amalgam of various techniques

• Secondary vertex tagging primary technique (SECVTX) – Identify tracks with large impact parameter

– Search for 2 or 3-track combinations with large L_{xy}

• Probability of tracks coming from primary vertex (JPB tag)

• Identify soft leptons in core of b-jet (SLT)

Tight and Loose B Tags

Employ concept of a “tight” and “loose” tag

• Tight ⇒ SECVTX tag

• Loose ⇒ Either SECVTX, JPB or SLT

• Allows trade-off between efficiency & bkg rejection

Use CDF Run I tagging efficiency

• Account for increase in SVX acceptance

• Use Run I mistag rates

Higgs Mass Resolution

Higgs mass resolution relies on jet reconstruction

• “Standard” CDF jet algorithm – gives M_{H} resolution ∼ 16 GeV/c^{2} at M_{H} = 110 GeV/c^{2}

• Track & shower-max data provides ∼ 20−30% improvement – Measured using photon-jet balancing

Assumed Improvement in Mass Resolution

Provides 10 − 20% improvement in H → bb̄ mass

This still requires work

• Topology dependence needs to be understood – Overlap with other objects in event – Combinatorial confusion

• Take into account other calorimetry effects

• Overall, promising development for Run II

Detection Strategy for Light SM Higgs

The most promising channels for a light SM Higgs are

p p → W^{+} H → l^{+} ν_{l} bb̄
p p → Z^{0} H → l^{+} l^{−} bb̄

• Identify e^{+} or μ^{+} with p_{T} > 20 GeV/c

• Require E̸_{T} > 20 GeV and reject dilepton events

• Only two jets with E_{T} > 25 GeV and > 10 GeV with |η| < 2 – Both b-tagged (one tight and one loose)

– Impose M_{bb} mass window ∼ ±15%

– No extra jets with E_{T} > 20 GeV and |η| < 2.4

Results in ϵ × B(H → bb̄) ∼ 2%

• Expected 8 fb observed for M_{H} = 90 GeV/c^{2} – Falls to 2 fb for M_{H} = 130 GeV/c^{2}

Background Rates for W H

All of the above come down to control of backgrounds

• Non-physical backgrounds estimated to be small

• Largest physical backgrounds are – QCD production of W bb̄

– Production of t t where two jets are missed

– Di-vector boson production

– Single top production (tb and tqb)

• Use HERWIG and PYTHIA, checked against Run I data Observed Cross Section (fb)

Process  M_{H} = 90  M_{H} = 110  M_{H} = 130
Signal  8.4  5.0  2.2
W bb̄  21.5  18.9  15.4
t t  9.6  12.2  13.0
V V  7.7  4.9  0.6
tb and tqb  9.4  11.8  12.7

Uncertainties in Signals and Backgrounds

Signal estimates suffer mostly from

• Uncertainties on b-tagging estimates

• Jet resolution estimates

• Uncertainties in production rates

• How large? SHW estimates are about 10-20% higher

What should you worry most about backgrounds?

• Rate of W bb̄ remains uncertain – Other physical backgrounds reasonably well understood

• Could W bb̄ be ×2 larger?

– Probably not...CDF estimate constrained by the b-tagging rates and kinematic data

– Also have other tools not used

∗ NN approach gets 50% lower background

• Remains a challenge in this analysis

Results for ZH Channel

ZH → l^{+} l^{−} bb̄ is straightforward

• Suffer about ×5 hit due to BR and lepton ID efficiency • Expect about 1/6 of W H observed rate

• Background rates are also approximately 1/6-1/7

ZH → ν_{l}ν̄_{l} bb̄ presents challenge

• Only have E̸_{T} signature from Z → νν̄

• Have two b-tagged jets reconstructing to H

• Creates a very “modest” signature – Issues with triggering

– Issues with beating down the backgrounds

Backgrounds to νν̄ bb̄ Final State

Physical backgrounds reasonably well understood

• QCD production of Zbb̄

• QCD production of W bb̄ (miss charged lepton)

• W Z and ZZ

• t t and single top

Observed cross section (fb)

Process  M_{H} = 90  M_{H} = 110  M_{H} = 130
Signal  3.6  2.7  0.9
Zbb̄  6.5  4.9  3.8
ZZ  5.4  3.4  0.0
t t  1.6  2.2  2.2
qq̄ → bb̄  ∼ 7  ∼ 6  ∼ 3

QCD bb̄ background difficult to estimate

• Depends crucially on E̸_{T} measurement

• Raw production cross section not well understood ⇒ Have to measure with data

Higher Mass Higgs Searches

Heavier Higgs (M_{H} > ∼ 130 GeV/c^{2}) rates lower

• H → W W^{∗} becomes dominant decay mode

• Production via gg → HX and qq̄′ → W H are important • Realistically, will require in excess of 15 fb^{−1}

This presents a number of decay channels

p p → H → W^{∗}W^{∗} → ll̄ν_{l}ν̄_{l}
p p → W^{+} H → W^{+} W^{∗}W^{∗} → l^{+} l^{+} qq̄′X

• These two represent the highest rate topologies

• Same-sign dileptons very clean

Signal and Background Estimates

Signal requirements for H → ll̄′ν_{l}ν̄_{l} are:

• Isolated leptons with p_{T} > 10 GeV/c and |η| < 1.5 • Require M_{ll′} < 78 GeV/c^{2} for e^{+} e^{−} and μ^{+} μ^{−}

– and M_{ll′} < 110 GeV/c^{2} for e^{+} μ^{−}

• Make decay angle cuts to reject τ^{+} τ^{−} events

• Reject events with high energy jets (eg., E_{T} > 95)

A more sophisticated likelihood analysis defines M_{C} ≡ √p_{T}^{2} (ll) + M^{2}(ll) + E̸_{T}

Heavy Higgs Signal and Background Rates

Estimates of signal and background rates

• Dominated by W^{+} W^{−} and t t

Observed Cross Section (fb)

Process  M_{H} = 140  M_{H} = 160  M_{H} = 180
Signal  2.6  1.5  1.0
Backgrounds  4.4  4.4  3.8
S/√B  2.1  3.9  2.7

Can estimate total integrated luminosity needed

• For either 3σ observation or 95% CL exclusion

• Include 10% systematic uncertainty

Associated Production of t tH

Recently, t tH has been reexamined at the Tevatron

Goldstein et al, hep-ph/0006311

• Analogous to bb̄H channel

• The t t gives efficient trigger and tag

Signal cross sections are small

• Estimate σ × B of t tH few fb

t tH Rates

Have to deal with very large t tX background

• t t + jj has rate of ∼ 1, 000 fb

• B-tagging is crucial to pull out signal – Require 2 jets, at least one b-tagged

• For M_{H} = 120 GeV/c^{2} and 15 fb^{−1} – 2-3 signal events and little background

Could yield a few “spectacular” events

Summary for SM Higgs Search

Results of the Workshop analysis can be combined:

• Use results of NN analysis

• Combine statistical power of both experiments

• Bands reflect 30% total additional “shift”

MSSM Higgs Search

Search for MSSM Higgs has to grapple with

• 5 Higgs bosons in minimal model – two CP -even scalars h and H – CP-odd scalar A

– pair of charged states H^{±} • A larger number of parameters

– tan β = v_{1}/v_{2} and M_{A}

Sketch “benchmark” measure of Tevatron sensitivity

• Look at two production models – qq̄ → V ϕ → V bb̄

– gg or qq̄ → bb̄ϕ → bb̄ bb̄

• Reinterpret limits from earlier analyses

R_{th}(X, Y) = \frac{σ_{X}(ϕ) B(ϕ → Y)}{σ_{X}(h_{SM}) B(h_{SM} → Y)}

Reinterpreting SM Search Results

Take results from SM W H → W bb̄ and ZH → Zbb̄

• Determine range of 95% exclusion and 5σ sensitivity

• Add 10% systematic uncertainty

R_{th}(qq̄′ → V ϕ; bb̄) ≡ (\frac{g_{V V ϕ}^{2}}{g_{V V H_{SM}}^{2}} ) [\frac{BR(ϕ → bb̄)}{BR(H_{SM} → bb̄)} ]

Can map these limits onto specific MSSM models

Looking at non-SM Channels

Final states that are unique to the MSSM Higgs

• SM production or decay rate is suppressed

• Results in a observable signature at Tevatron

• “Benchmark” example (for large tan β) is

gg or qq̄ → bb̄ϕ → bb̄ bb̄

Search for bb̄ bb̄ final states:

• Require 4 jets, E_{T} > 15 GeV and |η| < 2.1

• 1st and 2nd jet E_{T} > 40 − 75 GeV (M_{H} = 70 − 200 GeV/c^{2})

• Require 3 b-tagged jets

• Reconstruct ϕ → bb̄ signal

Higgs Reconstruction bb̄ Mode

Use a mass-dependent strategy to reconstruct H

• highest-E_{T} pair of jets for M_{H} > 120 GeV/c^{2}

• 2nd and 3rd highest-E_{T} jets for M_{H} ≤ 120 GeV/c^{2}

Estimated Signal and Background

Signal efficiency is relatively small

• Ranges from 0.4% to 0.5%

• Dominated by trigger requirements and b-tagging

Backgrounds are large (and somewhat uncertain)

• QCD background dominates (and largely irreducible)

• t t is next largest (about 1/5 of QCD)

Observed Cross Section (fb)

Process  M_{H} = 120  M_{H} = 140  M_{H} = 200
Signal  19  10  1.1
QCD  55 ± 12  34 ± 9  9 ± 3
t t  13 ± 3  11 ± 3  4 ± 1
S/√B  2.2  1.4  0.3

Summary of MSSM Higgs Mass Sensitivity

Can summarize these data

• 95% CL upper limits on tan β vs M_{H}

• Estimate 3σ and 5σ contours

Has extended reach for very large tan β values

• Large uncertainties in the QCD background

• b-tagging backgrounds are in “virgin” territory

Summary

Higgs search at the Tevatron shows clear promise

• Limited reach with 2 fb^{−1} (2002 goal)

• Data set of 15-20 fb^{−1} does have sensitivity – Could discover SM Higgs up to 115-125 GeV/c^{2}

– Sensitive to MSSM models over large parameter space

“Gold-plated” modes are

• p p → V H → V bb̄ for M_{H} < ∼ 130 GeV/c^{2}

• p p → H → V V^{∗} for M_{H} > ∼ 130 GeV/c^{2}

Many uncertainties in these predictions

• Based on Run I flavour-tagging experience

• Requires optimal kinematic reconstruction tools – Promising results on dijet mass reconstruction

– Have to be reproduced in Run II environment

– Resolution in E̸_{T} critical in some modes

• Will have to be resolved with the data